OpenStudy (anonymous):

Problem set 1, 1a-4 (b), is there a typo in the formula provided? The problem set has the numerator of second function on the right as f(x)+f(-x) whereas the solutions have it as f(x)-f(-x). I get that if the second function is an odd function then f(-x) = -f(x), but I'm getting myself tied in knots in terms of how that fits into the overall function provided. Thx!

OpenStudy (anonymous):

There is a typo in the question. It should read: \[f(x) = \frac{ f(x) + f(-x) }{ 2 } + \frac{ f(x) - f(-x)}{ 2 }\] If you add up the right side you will see that that this comes out to \[\frac{ f(x) + f(x) }{ 2 } = f(x)\] which is equal to the left side. The solution states the problem correctly. However, the solution also contains a typo. The last line of the solution should read: \[G(-x) = \frac{ f(-x) - f(-(-x)) }{ 2 } = \frac{ f(-x) - f(x) }{ 2 } = - G(x)\] I have encountered quite a few typos and discrepancies between the problems and solutions for this problem set.

OpenStudy (anonymous):

perfect - thanks!